2(-n+4)-5=2n+5 what does "n" equal??
Answer by sinadone(8)

2(-n+4)-5=2n+5
Distribute first.
2 * -n = -2n
2 * 4 = 8
So you have:
-2n+8-5
(add the 8 and 5)
-2n+3=2n+5
Move the 3 over to the right, making it negative.
-2n=2n+2
Then move the positive 2n to the left and add like terms:
-4n=2
Divide by -4
n=-1/2